# Classical and Quantum KMS States on Spin Lattice Systems

### v2

@@ 3 The semiclassical limit of the quantum KMS condition
In Section 3.1 we will recall the notion of classical and quantum KMS condition within the framework introduced in Section 2.1, cf. [13, 20]. In passing, we will prove a characterization of the classical KMS condition, cf. Lemma 3.2, which is inspired by an analogous characterization of the quantum KMS condition known with the name of Roepstorff-Araki-Sewell auto-correlation lower bound, cf. [13, Thm. 5.3.15]. Section 3.2 is devoted to the proof of Theorem 3.5. The latter is based on Lemma 3.2 together with a result on the semiclassical limit of the quantum derivation associated with the quantum KMS condition, cf. Lemma 3.4.
@@ 3.1 Classical and quantum KMS condition on $B_{\infty}^{\Gamma}$ and $B_{j}^{\Gamma}$
Within assumption (32) it can be shown that $\delta_{j}^{\Lambda}$ converges strongly on $\dot{B}_{j}^{\Gamma}$ as $\Lambda\uparrow\Gamma$ to a $C^{*}$-derivation $\boxed{\delta_{j}^{\Gamma}}\colon\dot{B}_{j}^{\Gamma}\to B_{j}^{\Gamma}$ which generates a strongly continuous one-parameter group $t\mapsto\tau^{\Gamma}_{t}$ on $B_{j}^{\Gamma}$, cf. [13, Thm. 6.2.4]. In particular, $\delta_{j}^{\Gamma}$ is explicitly given by
@@
where the series converges in $B_{j}^{\Gamma}$ for $|t|\leq\lambda/2\|\Phi_{j}\|_{\lambda}$, cf. [13, Thm. 6.2.4].
@@
    Remarkably, any weak*-limit point of the sequence $(\omega_{j}^{\beta,\Lambda})_{\Lambda\Subset\Gamma}$ of $(\beta,\delta_{j}^{\Lambda})$-KMS quantum states leads to a $(\beta,\delta_{j}^{\Gamma})$-KMS quantum state, cf. [13, Cor. 6.2.19]. The existence of these weak*-limit points follows from a standard Hahn-Banach and weak*-compactness argument. Similar considerations apply to the sequence $(\omega_{\infty}^{\beta,\Lambda})_{\Lambda\Subset\Gamma}$ of $(\beta,\delta_{\infty}^{\Lambda})$-KMS classical states, ensuring the existence of $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states.
@@ 3.2 Semi-classical limit of quantum KMS condition
The proof requires two main ingredients. The first one is a useful characterization of the $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical condition, cf. Lemma 3.2: The latter is inspired by an analogous characterization of $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states, cf. Remark 3.3. The second piece of information is a control of the semiclassical limit of the derivation $\delta_{j}^{\Gamma}$. Specifically, for any finite region $\Lambda\Subset\Gamma$, the DGR condition (3) implies that $h_{j}^{-1}\delta_{j}^{\Lambda}\circ Q_{j}^{\Lambda}-Q_{j}^{\Lambda}\circ\delta_{j}^{\Lambda}\underset{j\to\infty}{\longrightarrow}0$ strongly on $\dot{B}_{\infty}^{\Lambda}$. This is a key property which notably holds also for the lattice system on the entire $\Gamma$, cf. Lemma 3.4.
We begin with the characterization of the classical KMS condition. The following result is essentially the classical version of the Roepstorff-Araki-Sewell auto-correlation lower bound, cf. [13, Thm. 5.3.15].
@@
To begin with we observe that the classical auto-correlation lower bound implies in particular that $-i\omega(\mathfrak{a}^{*}\delta(\mathfrak{a}))\geq 0$ for all $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$: In particular $-i\omega(\mathfrak{a}^{*}\delta(\mathfrak{a}))\in\mathbb{R}$ for all $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$. This implies that $\omega(\delta(\mathfrak{a}))=0$ for all $\mathfrak{a}\in\dot{\mathfrak{A}}$: The proof is a classical counterpart of [13, Lem. 5.3.16] and will be reviewed for the sake of clarity. In particular for $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$ we find
@@
The classical auto-correlation lower bound (39) is a classical analogue of the quantum auto-correlation lower bound, cf. [13, Thm. 5.3.15]. To state the latter let $\mathfrak{A}$ be a non-commutative $C^{*}$-algebra and let $\tau$ be a strongly continuous one-parameter group of *-automorphisms on $\mathfrak{A}$ with infinitesimal generator $\delta$. Then $\omega\in S(\mathfrak{A})$ is a $(\beta,\delta)$-KMS quantum state if and only if
@@
where $\otimes_{\varepsilon}$ denotes the injective tensor product of Banach spaces, cf. [47, §3.2]. Let $T_{j}\colon\dot{B}_{\infty}^{\Gamma}\to\dot{B}_{j}^{\Gamma}$ be the linear operator defined by
@@
We then consider the inductive tensor product of $\hat{T}_{j}^{\Lambda}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}$ with $Q_{j}^{\Lambda^{c}}\colon B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Lambda^{c}}$, cf. [47, Prop. 3.2]. The latter is the unique bounded linear map
@@
for all $b_{\Lambda}\in C^{1}(\mathbb{S}^{2}_{\Lambda})$ and $b_{\Lambda^{c}}\in B_{\infty}^{\Lambda^{c}}$. Notice that $\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}$ is a $C^{*}$-algebra once completed with its unique $C^{*}$-cross norm. The uniqueness of such norm is due to the fact that $B_{j}^{\Lambda^{c}}$ is the $C^{*}$-inductive limit of finite dimensional, hence nuclear, $C^{*}$-algebras, therefore, it is nuclear as well, cf. [10, §II.9.4.5]. We recall that
@@
4 Absence of CPTs implies absence of QPTs
@@
To this avail we will consider [13, Prop. 6.2.45], which provides a mild sufficient condition on a family of quantum potential $\Phi_{j}=\{\Phi_{j,\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness of the corresponding $(\beta,\delta_{j}^{\Gamma})$-KMS quantum state on $B_{j}^{\Gamma}$ for all $\beta\leq\beta_{j}$, where $\beta_{j}$ is called quantum critical inverse temperature. This result does not suffices to our purposes because: (a) there is no clear classical counterpart for the assumption on the family $\Phi_{j}$; (b) the estimate for the critical inverse temperature $\beta_{j}$ provided in [13, Prop. 6.2.45] is not uniform in $j\in\mathbb{Z}_{+}/2$, spoiling a comparison with the classical critical inverse temperature —cf. Remark 4.9.
To encompass these issues we revised the proof of [13, Prop. 6.2.45] in other to produce a new result which is uniform in $j\in\mathbb{Z}_{+}/2$. In particular, in Section 4.1 we prove a classical analogous of [13, Prop. 6.2.45], cf. Theorem 4.1. The latter result provides a sufficient condition, cf. (44), on the family of potential $\varphi=\{\varphi_{\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness for the corresponding $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical state on $B_{\infty}^{\Gamma}$ provided $\beta\leq\beta_{\infty}$, $\beta_{\infty}$ being the classical critical inverse temperature. Condition (44) is essentially a classical version of the one required in [13], cf. (72), and essentially boils down to an estimate on the sup-norm of the derivatives of the potentials $\varphi_{\Lambda}$, $\Lambda\Subset\Gamma$.
Next, we provide a uniqueness result, cf. Theorem 4.7, for $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states on $B_{j}^{\Gamma}$ associated with the family of quantum potentials $\Phi_{j,\Lambda}=Q_{j}^{\Gamma}(\varphi_{\Lambda})$ obtained by Berezin quantization of the classical potential. The novelty of this result relies on the bound on the critical inverse temperature, which is now uniform in $j\in\mathbb{Z}_{+}/2$, allowing for a better comparison with the classical critical inverse temperature. Finally, it is shown that the assumptions of Theorem 4.7 implies those of Theorem 4.1, cf. Remark 4.9. This shows in this particular model that absence of classical phase transitions and quantum phase transitions coexists for high enough temperatures.
@@ 4.1 Uniqueness result for classical KMS states
The goal of this section is to prove a uniqueness result for $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states, cf. Theorem 4.1. The latter identifies a sufficient condition on the classical interaction potentials $\varphi=\{\varphi_{\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness of $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states on $B_{\infty}^{\Gamma}$ for sufficiently low $\beta$, that is, at sufficiently high temperature. This result is inspired from [13, Prop. 6.2.45] and adapted to the classical setting. In the forthcoming sections we will present an analogous uniqueness result in the quantum setting, cf. Theorem 4.7.
@@
    Lemma 4.4 is inspired by [22, Prop. 5], where Equation (52) has been proved in the context of finite dimensional Poisson manifolds. Equation 52 can be understood as an adaptation to the present setting.
@@
To begin with, we choose an arbitrary but fixed bijection $\Gamma\simeq\mathbb{Z}$ and induce an ordering on $\Gamma$ based on such map. Thus, for any $\Lambda\Subset\Gamma$ we may set $x:=\min_{y\in\Lambda}y$ where the minimum is taken with respect to the chosen ordering.
@@ 4.2 Uniqueness result for quantum KMS state
The goal of this section is to prove a uniqueness result for $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states on $B_{j}^{\Gamma}$, cf. Theorem 4.7 The latter applies under hypothesis very similar to those of Theorem 4.1. In fact, Theorems 4.1, 4.7 will imply that, under suitably mild assumptions, for high enough temperatures there is absence of both classical and quantum phase transitions, cf. Remark 4.9. The proof of Theorem 4.7 is inspired by [13, Prop. 6.2.45], see also [28], although it requires a different argument to ensure a uniform bound on $j\in\mathbb{Z}_{+}/2$.
@@
    where we used the explicit expression obtained in Example 2.1.
@@
    For later convenience we also discuss a quantum version of Equation (17). Indeed, once again thanks to Equation (23), an analogous identity holds for the $\mathcal{Y}$’s. This can be either argued by observing that the $\mathcal{Y}_{j|\ell,m}$ are spherical tensors of order $\ell$ with respect to the representation $\tilde{D}^{(j)}$, cf. [42, §3.11], or by direct inspection. In more details, let $\ell_{1},\ell_{2}\in\mathbb{Z}_{+}$ and $m_{1},m_{2}\in\mathbb{Z}$ with $m_{k}\in[-\ell_{k},\ell_{k}]$, $k\in\{1,2\}$. Equation (65) leads to
@@
    where we used a few properties of the Clebsch-Gordan coefficients and the definition of the $6j$-symbols, cf. [34, §8.7 and §9.1]. Notice that the appearance of the Clebsch-Gordan coefficient $\textsc{cg}_{\ell_{1},m_{1};\ell_{2},m_{2}}^{\ell,m}$ ensures that the coefficients vanish unless $|\ell_{1}-\ell_{2}|\leq\ell\leq\ell_{1}+\ell_{2}$. Moreover, since it is known that the matrix
@@
Finally we apply the following estimate, cf. [13, Prop. 6.2.45]: If $\alpha_{X}\in\mathbb{R}_{+}$ for all $X\Subset\Gamma$ then for all $S\Subset\Gamma$
@@
    It is worth to compare Theorem 4.7 with [13, Prop. 6.2.45]. The latter provide a sufficient condition for the uniqueness of $(\beta,\delta_{j}^{\Gamma})$-KMS for fixed $j\in\mathbb{Z}_{+}/2$ which is similar in spirit to (61) —in fact, Theorem 4.7 has been inspired by this latter result— namely
@@
    Condition (72) is stronger than (61) because it only uses the $B_{j}^{X}$-norm of $Q_{j}^{X}(\varphi_{X})$. However, it is not uniform of $j\in\mathbb{Z}_{+}/2$, in particular, it requires a faster and faster decay behaviour of the potential $\{\varphi_{\Lambda}\}_{\Lambda\Subset\mathbb{Z}^{d}}$ as $j\to\infty$. Moreover, the critical quantum inverse temperature $\beta_{\textsc{br}}(j,\lambda)$ predicted by [13, Prop. 6.2.45] vanishes as $j\to\infty$. Finally, there is no classical version of condition (72). For all these reasons [13, Prop. 6.2.45] is not suitable for the comparison with the classical setting we are interested in.

### v3

@@ 3 The semiclassical limit of the quantum KMS condition
In Section 3.1 we will recall the notion of classical and quantum KMS condition within the framework introduced in Section 2.1, cf. [15, 22]. In passing, we will prove a characterization of the classical KMS condition, cf. Lemma 3.2, which is inspired by an analogous characterization of the quantum KMS condition known with the name of Roepstorff-Araki-Sewell auto-correlation lower bound, cf. [15, Thm. 5.3.15]. Section 3.2 is devoted to the proof of Theorem 3.5. The latter is based on Lemma 3.2 together with a result on the semiclassical limit of the quantum derivation associated with the quantum KMS condition, cf. Lemma 3.4.
@@ 3.1 Classical and quantum KMS condition on $B_{\infty}^{\Gamma}$ and $B_{j}^{\Gamma}$
Within assumption (32) it can be shown that $\delta_{j}^{\Lambda}$ converges strongly on $\dot{B}_{j}^{\Gamma}$ as $\Lambda\uparrow\Gamma$ to a $C^{*}$-derivation $\boxed{\delta_{j}^{\Gamma}}\colon\dot{B}_{j}^{\Gamma}\to B_{j}^{\Gamma}$ which generates a strongly continuous one-parameter group $t\mapsto\tau^{\Gamma}_{t}$ on $B_{j}^{\Gamma}$, cf. [15, Thm. 6.2.4]. In particular, $\delta_{j}^{\Gamma}$ is explicitly given by
@@
where the series converges in $B_{j}^{\Gamma}$ for $|t|\leq\lambda/2\|\Phi_{j}\|_{\lambda}$, cf. [15, Thm. 6.2.4].
@@
    Remarkably, any weak*-limit point of the sequence $(\omega_{j}^{\beta,\Lambda})_{\Lambda\Subset\Gamma}$ of $(\beta,\delta_{j}^{\Lambda})$-KMS quantum states leads to a $(\beta,\delta_{j}^{\Gamma})$-KMS quantum state, cf. [15, Cor. 6.2.19]. The existence of these weak*-limit points follows from a standard Hahn-Banach and weak*-compactness argument. Similar considerations apply to the sequence $(\omega_{\infty}^{\beta,\Lambda})_{\Lambda\Subset\Gamma}$ of $(\beta,\delta_{\infty}^{\Lambda})$-KMS classical states, ensuring the existence of $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states.
@@ 3.2 Semi-classical limit of quantum KMS condition
The proof requires two main ingredients. The first one is a useful characterization of the $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical condition, cf. Lemma 3.2: The latter is inspired by an analogous characterization of $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states, cf. Remark 3.3. The second piece of information is a control of the semiclassical limit of the derivation $\delta_{j}^{\Gamma}$. Specifically, for any finite region $\Lambda\Subset\Gamma$, the DGR condition (3) implies that $h_{j}^{-1}\delta_{j}^{\Lambda}\circ Q_{j}^{\Lambda}-Q_{j}^{\Lambda}\circ\delta_{\infty}^{\Lambda}\underset{j\to\infty}{\longrightarrow}0$ strongly on $\dot{B}_{\infty}^{\Lambda}$. This is a key property which notably holds also for the lattice system on the entire $\Gamma$, cf. Lemma 3.4.
We begin with the characterization of the classical KMS condition. The following result is essentially the classical version of the Roepstorff-Araki-Sewell auto-correlation lower bound, cf. [15, Thm. 5.3.15].
@@
To begin with we observe that the classical auto-correlation lower bound implies in particular that $-i\omega(\mathfrak{a}^{*}\delta(\mathfrak{a}))\geq 0$ for all $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$: In particular $-i\omega(\mathfrak{a}^{*}\delta(\mathfrak{a}))\in\mathbb{R}$ for all $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$. This implies that $\omega(\delta(\mathfrak{a}))=0$ for all $\mathfrak{a}\in\dot{\mathfrak{A}}$: The proof is a classical counterpart of [15, Lem. 5.3.16] and will be reviewed for the sake of clarity. In particular for $\mathfrak{a}=\mathfrak{a}^{*}\in\dot{\mathfrak{A}}$ we find
@@
The classical auto-correlation lower bound (39) is a classical analogue of the quantum auto-correlation lower bound, cf. [15, Thm. 5.3.15]. To state the latter let $\mathfrak{A}$ be a non-commutative $C^{*}$-algebra and let $\tau$ be a strongly continuous one-parameter group of *-automorphisms on $\mathfrak{A}$ with infinitesimal generator $\delta$. Then $\omega\in S(\mathfrak{A})$ is a $(\beta,\delta)$-KMS quantum state if and only if
@@
where $\otimes_{\varepsilon}$ denotes the injective tensor product of Banach spaces, cf. [51, §3.2]. Let $T_{j}\colon\dot{B}_{\infty}^{\Gamma}\to\dot{B}_{j}^{\Gamma}$ be the linear operator defined by
@@
We then consider the inductive tensor product of $\hat{T}_{j}^{\Lambda}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}$ with $Q_{j}^{\Lambda^{c}}\colon B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Lambda^{c}}$, cf. [51, Prop. 3.2]. The latter is the unique bounded linear map
@@
for all $b_{\Lambda}\in C^{1}(\mathbb{S}^{2}_{\Lambda})$ and $b_{\Lambda^{c}}\in B_{\infty}^{\Lambda^{c}}$. Notice that $\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}$ is a $C^{*}$-algebra once completed with its unique $C^{*}$-cross norm. The uniqueness of such norm is due to the fact that $B_{j}^{\Lambda^{c}}$ is the $C^{*}$-inductive limit of finite dimensional, hence nuclear, $C^{*}$-algebras, therefore, it is nuclear as well, cf. [12, §II.9.4.5]. We recall that
@@
4 Common absence of CPTs and of QPTs at high temperatures
@@
To this avail we will consider [15, Prop. 6.2.45], which provides a mild sufficient condition on a family of quantum potential $\Phi_{j}=\{\Phi_{j,\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness of the corresponding $(\beta,\delta_{j}^{\Gamma})$-KMS quantum state on $B_{j}^{\Gamma}$ for all $\beta\leq\beta_{j}$, where $\beta_{j}$ is called quantum critical inverse temperature. This result does not suffices to our purposes because: (a) there is no clear classical counterpart for the assumption on the family $\Phi_{j}$; (b) the estimate for the critical inverse temperature $\beta_{j}$ provided in [15, Prop. 6.2.45] is not uniform in $j\in\mathbb{Z}_{+}/2$, spoiling a comparison with the classical critical inverse temperature —cf. Remark 4.9.
To encompass these issues we revised the proof of [15, Prop. 6.2.45] in order to produce a new result which is uniform in $j\in\mathbb{Z}_{+}/2$. In particular, in Section 4.1 we prove a classical analogous of [15, Prop. 6.2.45], cf. Theorem 4.1. The latter result provides a sufficient condition, cf. (44), on the family of potential $\varphi=\{\varphi_{\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness for the corresponding $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical state on $B_{\infty}^{\Gamma}$ provided $\beta\leq\beta_{\infty}$, $\beta_{\infty}$ being the classical critical inverse temperature. Condition (44) is essentially a classical version of the one required in [15], cf. (72), and essentially boils down to an estimate on the sup-norm of the derivatives of the potentials $\varphi_{\Lambda}$, $\Lambda\Subset\Gamma$.
Next, we provide a uniqueness result, cf. Theorem 4.7, for $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states on $B_{j}^{\Gamma}$ associated with the family of quantum potentials $\Phi_{j,\Lambda}=Q_{j}^{\Gamma}(\varphi_{\Lambda})$ obtained by Berezin quantization of the classical potential. The novelty of this result relies on the bound on the critical inverse temperature, which is now uniform in $j\in\mathbb{Z}_{+}/2$, allowing for a better comparison with the classical critical inverse temperature. Finally, it is shown that the assumptions of Theorem 4.7 implies those of Theorem 4.1, cf. Remark 4.9. For this class of models, this establishes a sufficient criterion for the absence of both classical and quantum phase transitions.
@@ 4.1 Uniqueness result for classical KMS states
The goal of this section is to prove a uniqueness result for $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states, cf. Theorem 4.1. The latter identifies a sufficient condition on the classical interaction potentials $\varphi=\{\varphi_{\Lambda}\}_{\Lambda\Subset\Gamma}$ which ensures uniqueness of $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical states on $B_{\infty}^{\Gamma}$ for sufficiently low $\beta$, that is, at sufficiently high temperature. This result is inspired from [15, Prop. 6.2.45] and adapted to the classical setting. In the forthcoming sections we will present an analogous uniqueness result in the quantum setting, cf. Theorem 4.7.
@@
    Lemma 4.4 is inspired by [24, Prop. 5], where Equation (52) has been proved in the context of finite dimensional Poisson manifolds. Equation 52 can be understood as an adaptation to the present setting.
@@
To begin with, we choose an arbitrary but fixed bijection $\Gamma\simeq\mathbb{Z}$ and induce an ordering on $\Gamma$ based on such map. Thus, for any $\Lambda\Subset\Gamma$ we may set $x:=\min_{y\in\Lambda}y$ where the minimum is taken with respect to the chosen ordering. The choice of the ordering is only made to select a distinguished point $x\in\Lambda$ for any $\Lambda\Subset\Gamma$.
@@ 4.2 Uniqueness result for quantum KMS state
The goal of this section is to prove a uniqueness result for $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states on $B_{j}^{\Gamma}$, cf. Theorem 4.7 The latter applies under hypothesis very similar to those of Theorem 4.1. In fact, Theorems 4.1, 4.7 will imply that, under suitably mild assumptions, for high enough temperatures there is absence of both classical and quantum phase transitions, cf. Remark 4.9. The proof of Theorem 4.7 is inspired by [15, Prop. 6.2.45], see also [30], although it requires a different argument to ensure a uniform bound on $j\in\mathbb{Z}_{+}/2$.
@@
    where we used the explicit expression obtained in Example 2.1. It is worth to mention that further properties of the matrices $\mathcal{Y}_{j|\ell,m}$ have been studied in [34].
@@
    For later convenience we also discuss a quantum version of Equation (17). Indeed, once again thanks to Equation (23), an analogous identity holds for the $\mathcal{Y}$’s. This can be either argued by observing that the $\mathcal{Y}_{j|\ell,m}$ are spherical tensors of order $\ell$ with respect to the representation $\tilde{D}^{(j)}$, cf. [45, §3.11], or by direct inspection. In more details, let $\ell_{1},\ell_{2}\in\mathbb{Z}_{+}$ and $m_{1},m_{2}\in\mathbb{Z}$ with $m_{k}\in[-\ell_{k},\ell_{k}]$, $k\in\{1,2\}$. Equation (65) leads to
@@
    where we used a few properties of the Clebsch-Gordan coefficients and the definition of the $6j$-symbols, cf. [37, §8.7 and §9.1]. Notice that the appearance of the Clebsch-Gordan coefficient $\textsc{cg}_{\ell_{1},m_{1};\ell_{2},m_{2}}^{\ell,m}$ ensures that the coefficients vanish unless $|\ell_{1}-\ell_{2}|\leq\ell\leq\ell_{1}+\ell_{2}$. Moreover, since it is known that the matrix
@@
Finally we apply the following estimate, cf. [15, Prop. 6.2.45]: If $\alpha_{X}\in\mathbb{R}_{+}$ for all $X\Subset\Gamma$ then for all $S\Subset\Gamma$
@@
    It is worth to compare Theorem 4.7 with [15, Prop. 6.2.45]. The latter provide a sufficient condition for the uniqueness of $(\beta,\delta_{j}^{\Gamma})$-KMS for fixed $j\in\mathbb{Z}_{+}/2$ which is similar in spirit to (61) —in fact, Theorem 4.7 has been inspired by this latter result— namely
@@
    Condition (72) is stronger than (61) because it only uses the $B_{j}^{X}$-norm of $Q_{j}^{X}(\varphi_{X})$. However, it is not uniform of $j\in\mathbb{Z}_{+}/2$, in particular, it requires a faster and faster decay behaviour of the potential $\{\varphi_{\Lambda}\}_{\Lambda\Subset\mathbb{Z}^{d}}$ as $j\to\infty$. Moreover, the critical quantum inverse temperature $\beta_{\textsc{br}}(j,\lambda)$ predicted by [15, Prop. 6.2.45] vanishes as $j\to\infty$. Finally, there is no classical version of condition (72). For all these reasons [15, Prop. 6.2.45] is not suitable for the comparison with the classical setting we are interested in.